Jieen:

Chords DB and MN intersect at the center of circle MBND . Which is the measure of minor arc DM

Narad:

The length of the arc will depend on the angle between DB and MN. If the radius of the circle is r, then \[0 \le "length of arc" \le \pi*r\]